OpenStudy (anonymous):

How many ways can 4 students be chosen from a class of 12 to represent their class at a banquet? 495 990 5,940 11,880

OpenStudy (anonymous):

you are asked for "12 choose 4" written as \[\binom{12}{4}\] or \[_{12}C_{4}\] or even \[^{12}C_4\] and computed via \[\binom{12}{4}=\frac{\overbrace{12\times 11\times 10\times 9}^{\text{ 4 terms}}}{4\times 3\times 2}\] cancel first, multiply last

OpenStudy (anonymous):

or just cheat http://www.wolframalpha.com/input/?i=12+choose+4